# The Junction of Immersive Analytics and Virtual Reconstructions – A Case Study on the Mausoleum of Emperor Maxentius thanks: DFG, project ID 251654672 – TRR 161

### v2

@@ V Discussion
With this work, we present three digital reconstructions of the Maxentius Mausoleum in Rome and utilize them to assess the value of IA for Virtual Archaeology. The reconstructions we propose for the Maxentius Mausoleum adhere to the typology of late antique imperial mausolea and are based on archaeological evidence and scholarly interpretations. Such scientific evidence provides us with a sound foundation for our discussion in which we want to advocate for the use of immersive technologies more prominently for research purposes. While there exist works that point out the usefulness of immersive environments in Virtual Archaeology [29], they tend to focus on dissemination or reconstruction and not on how to improve the perception of a specific aspect, like the comparison of hypotheses.
@@
Using the three reconstruction models, we created immersive reconstructions in the form of an interactive VR (iVR) environment. This iVR environment serves as proof of concept of the potential of IA in archaeological analysis. We integrated a selection of IA techniques to demonstrate a promising approach to accessing archaeological data that goes beyond representation, offering new ways to interact with and comprehend complex spatio-temporal data. Our iVR environment allows users to manipulate lighting conditions, toggle between showing a color coding for the certainty of single parts of the reconstruction and plain grey, place spatial markers and move around with a continuous movement metaphor. They use their hands to change the lighting and gain haptic feedback when using buttons inside the virtual space. We chose those interactions to align with IA research suggesting that embodied tools, among other concepts, can improve data comprehension [11]. We chose VR because the ability to interact with the reconstruction in an immersive environment supports a deeper understanding of spatial relationships. This is consistent with findings on the benefits of stereoscopic 3D and head-tracking for spatial judgment accuracy [28]. The color coding we use builds upon established practices in Virtual Archaeology to represent reliability in reconstructions [9].
@@
As the feature space of IA topics and Virtual Archaeology is vast, we decided against integrating an exhaustive number of potential features into our software. Our objective was not to provide comprehensive coverage but to demonstrate a number of possibilities as an example. The interaction possibilities we offer allow observers to gain a general impression of the reconstruction alternatives by switching through them on demand. Furthermore, observers can gain an intuition of the relative proportions of known parts of the mausoleum versus those that are not known but can be derived from similar architecture. Through the miniature, users can see how the light source’s position changes relative to the reconstruction.
The general implications of this work are significant for both archaeological methodology and public engagement. It emphasizes a number of ways to interact with and analyze archaeological data in terms of representation and metadata and argues where future virtual archaeology projects could draw inspiration from. The ability to explore spatial relationships, lighting conditions, and levels of certainty in an immersive environment with embedded analysis information could lead to new research questions.
@@
Due to the character of this work, there are a number of limitations to consider. As for the models we created, as is usual in the domain, they also have parts that are more interpretative than others. The current implementation is focused on a single site and a specific set of interaction techniques. Although multi-modality is important, only haptic feedback and a visual representation of the models are available. Additionally, the use of VR hardware limits the accessibility for some users. While we do offer simple controls that allow the software without VR hardware, the immersion factor that comes with VR is lost. Also, the considerations we make are a number of potential intersections of the two fields without taking deep considerations of how the design space of the different methods, like interaction design, would have to be applied.
@@
Thus, future work should include a variety of aspects. For one, a unified framework for creating immersive experiences would pose an asset. Such a framework should cover the inclusion and on-demand availability of historical background inside an immersive environment or the sonification of information, both of which have benefits. It may also be helpful to have ways of adjusting the light conditions tailored to the object’s geolocation and a chosen time and date.
@@
However, not only software is required. A relevant direction is to assess specific and established workflows from IA and how they can be transferred and used in Virtual Archaeology. It could also be valuable to compare the effectiveness of immersive environments with traditional methods of archaeological data analysis. Lastly, it would be interesting to see the effect of new metaphors for representing inherent information, such as levels of certainty, volumes of objects, sizes, usage, or chronological embedding.
@@
Therefore, there are a great number of opportunities in the future in this intersection, which should be discovered and evaluated together with domain experts to gain further insights into the usage of the Maxentius Mausoleum specifically and cultural heritage objects in general.
@@ VI Conclusion
In this work, we present three reconstruction hypotheses of the Maxentius Mausoleum and the literature and typology that inspired them. In addition, we have created an immersive environment in which those reconstructions can be viewed and explored. The environment comes with a small number of integrated interaction and navigation possibilities to provide not only the models but also the means to explore them. Our second contribution is the exploration of the potential for Virtual Archaeology that Immersive Analytics as a research direction yields. We advocate the shift from ever-better representations, recording, and storage of cultural heritage items to tailored immersive experiences that improve accessibility and insight into existing data. The direction is to look at how virtual archaeology studies can be optimized by identifying the specific requirements of a given problem, like comparing different monument reconstruction interpretations and finding out how to support that. For argumentation, we outline the possibilities of engaging representations of reconstructions and interactive representations of hypotheses. We exemplify the concept of integrating IA with our current research on the Maxentius Mausoleum, where these concepts will facilitate the analysis of the hypotheses regarding its construction and use.
In our future research, we want to look into how the workflows from IA can be integrated into digital reconstruction. We like to combine existing modeling workflows for cultural heritage objects and advancements in data representation in immersive environments. We will continue to utilize the Maxentius Mausoleum for that and branch out to other objects, where appropriate. Overall, we present an exciting research direction that offers much potential for enhancing users’ experience of virtual archaeology environments.

### v3

@@ V Discussion
With this work, we present three digital reconstructions of the Maxentius Mausoleum in Rome and utilize them to assess the value of IA for virtual archaeology. The reconstructions we propose for the Maxentius Mausoleum adhere to the typology of late antique imperial mausolea, based on archaeological evidence and scholarly interpretations. Such scientific evidence provides a sound foundation for our discussion, in which we aim to advocate for the more prominent use of immersive technologies to support archaeological research. While there exist works that highlight the usefulness of immersive environments in virtual archaeology [29], they tend to focus on dissemination or reconstruction rather than improving the perception of a specific aspect, such as the comparison of hypotheses.
From a historical and archaeological perspective, the immersive reconstruction of Maxentius’ mausoleum provides important insights into the changing relationship between light and space, a topic much debated in discussions on the development of Late Antique architecture. As an outstanding example of a new type of two-storeyed mausolea combining a dark crypt with light-flooded upper floors, it adds a completely new dimension of including light to funeral architecture that was taken up and further developed in the fourth century by church-attached imperial mausolea [Hesberg1992]. The graded lightning seen here became an important feature of late antique interiors, extending from funeral to palatial [gunter1968wand] and ecclesiastical [deBlaauw2010] architecture. An immersive reconstruction of Maxentius’ mausoleum complex will explore this dimension of light and lightning in a systematical manner and, at the same time, offer a promising starting point to take the general discussion of this important feature of late antique architecture to a new level.
@@
Thus, using the three reconstruction models, we created immersive reconstructions in the form of an interactive VR (iVR) environment. This iVR environment serves as proof of concept of the potential of IA in archaeological analysis. We integrated a selection of IA techniques to demonstrate a promising approach to accessing archaeological data that goes beyond representation. Our environment enables users to manipulate lighting conditions, toggle between color-coded representations of reconstruction certainty and grey, place spatial markers, and navigate with a continuous movement metaphor. They use their hands to adjust the lighting and gain haptic feedback when using buttons within the virtual space. We chose those interactions to align with IA research suggesting that embodied tools, among other concepts, can improve data comprehension [11]. We chose VR because it enables interaction with the reconstruction in an immersive environment, supporting a deeper understanding of spatial relationships. This finding is consistent with findings on the benefits of stereoscopic 3D and head-tracking for spatial judgment accuracy [28]. The color coding we use builds upon established practices in virtual archaeology to represent reliability in reconstructions [9].
@@
As the feature space of IA topics and virtual archaeology is vast, we decided against integrating an exhaustive number of potential features into our software. Our objective was not to provide comprehensive coverage but to demonstrate a number of possibilities as an example. The interaction possibilities we offer allow observers to gain a general impression of the reconstruction alternatives by switching between them on demand. Furthermore, observers can gain an intuition of the relative proportions of known parts of the mausoleum versus those that are not known but can be derived from similar architecture. Through the miniature, users can see how the position of the light source changes relative to the reconstruction.
@@
Due to the nature of this work, several limitations should be considered. As for the models we created, as is usual in the domain, they also have parts that are more interpretative than others. The current implementation is focused on a single site and a specific set of interaction techniques. Although multi-modality is important, only haptic feedback and a visual representation of the models are available. Also, the use of VR hardware limits accessibility for some users. While we do offer simple controls that allow the software to be used without VR hardware, the immersion factor that comes with VR is lost. Additionally, the considerations we make involve several potential intersections of the two fields, without taking into account the in-depth considerations of how the design space of different methods, such as interaction design, would need to be applied.
@@
For future work, a unified framework for creating immersive experiences would pose an asset. Such a framework should encompass the inclusion and on-demand availability of historical context within an immersive environment. This could happen in the form of integrated analysis tools, such as classifiers for identifying anachronistic components within a digitization. The sonification of information may also be of interest. It may also be helpful to have ways of adjusting the light conditions tailored to the object’s geolocation and a chosen time, date, and seasonal characteristics, which was out of the scope of this work.
@@
However, not only software is required. A relevant direction is to assess specific and established workflows from IA and how they can be transferred and used in virtual archaeology. It could also be valuable to compare the effectiveness of immersive environments with traditional methods of archaeological data analysis by evaluating the performance of experts on related tasks, such as identifying archaeological units. Lastly, it would be interesting to see the effect of new metaphors for representing inherent information, such as levels of certainty, volumes of objects, sizes, usage, or chronological embedding.
@@
Therefore, there are numerous opportunities in the future at this intersection, which should be explored and evaluated together with domain experts to gain further insights into the use of the Maxentius Mausoleum, specifically, and cultural heritage objects in general.
@@ VI Conclusion
In this work, we present three reconstruction hypotheses of the Maxentius Mausoleum, along with the literature and typology that inspired them. We created 3D models of those reconstruction hypotheses and provide them openly. In addition, we have created an immersive environment in which these reconstructions can be viewed and explored through a small number of integrated interaction and navigation options, either as VR or desktop applications. Furthermore, we explore the potential for virtual archaeology that immersive analytics as a research direction yields. We advocate the shift from ever-better representations, recording, and storage of cultural heritage items to tailored immersive experiences that improve accessibility and insight into existing data. The goal is to look at how virtual archaeology studies can be optimized by identifying the specific requirements of a given problem, such as comparing different monument reconstruction interpretations and investigating support options. For argumentation, we outline the possibilities of engaging representations of reconstructions and interactive representations of hypotheses. We exemplify the concept of integrating IA with our current research on the Maxentius Mausoleum, where these concepts will facilitate the analysis of hypotheses regarding its construction, use, and the relationship between light and space within it. Overall, we present an exciting research direction that offers much potential for enhancing users’ experience of virtual archaeology environments.